# CROP: Circuit Retrieval and Optimization with Parameter Guidance using LLMs

### v1

@@ I Introduction
To address this challenge, a promising research direction leverages circuit characteristics to enhance parameter tuning. This circuit-aware approach incorporates prior knowledge of design structure and functionality to more efficiently constrain the optimization process. VLSI design flows implementing such circuit-aware parameter tuning have demonstrated significant improvements in search efficiency. [13] used knowledge provided from circuit designers to guide the genetic algorithm and improve search efficiency. [11] leveraged a graph neural network and a transformer to help RL agent perform fast parameter exploration on unseen designs. [4] used random embedding to convert the parameters to a lower-dimensional latent space, which can be more efficiently explored based on design similarity. [10] exploited archival design data to improve the warm-up efficiency of the search engine. Despite these advances, significant challenges remain in effectively transferring knowledge in a way that better captures the correlation between pre-existing archival design data and the design currently being optimized. This is due to insufficient research on the effective encoding of the semantic representations of circuit designs. An approach that can understand circuit semantics and leverage knowledge from similar designs could dramatically reduce the number of iterations required to find optimal parameters.
Inspired by recent advances in LLMs and their application in EDA and chip design [14, 15, 16], we explore LLMs as a promising candidate for a fully-automated circuit-aware parameter search engine. Unlike traditional block-box optimization tools, LLMs can be conditioned using rich descriptive prompts that incorporate the QoR target, the functionality of each tool option, and even high-level descriptions of the circuit under optimization. This allows LLMs to emulate the intuition and practice of experienced engineers, making context-aware adjustments to EDA tool configurations, design constraints, and metric-oriented strategies. Furthermore, RAG emerges as a systematic approach that augments general-purpose LLMs with a domain-specific knowledge database, enabling more informed parameter optimization decisions based on historical design data and established optimization patterns. The integration of RAG with LLMs creates a parameter tuning framework that leverages contextual reasoning with circuit-specific prior knowledge, enabling parameter optimization that adapts to unique design characteristics across diverse technology nodes.
@@ II Preliminaries
Figure 1: Overview of CROP. In stage 1, CROP constructs a design database. CROP first repeats the LLM-based design analysis to generate the design embedding. CROP selects the top-$k$ parameter sets and their QoR results from existing optimization data to compile the parameter guidance for pre-existing circuits. In stage 2, CROP performs LLM-based analysis on the design of user query and retrieves the parameter guidance of the design from the database with highest similarity based on maximum inner product search (MIPS). The retrieved parameter guidance is used to constrain the follow-up LLM-based parameter search engine that optimizes the QoR for the input design.
@@ II-A Transformer-based Language Models
The advent of transformer-based language models has revolutionized the field of natural language processing (NLP) resulting in significant advancements in numerous tasks such as question answering, text generation, and machine translation. Introduced in 2017 by [17] the transformer architecture replaced recurrence with multihead self-attention, a mechanism that allowed more efficient training and better performance in sequence modeling tasks. Subsequent developments of larger models such as GPT3 [18] and LLaMA3 [19], have shown strong zero-shot and few-shot generalization skills when trained on massive corpora. These models are typically pre-trained for next-token generation and later finetuned for downstream tasks such as instruction following or question answering. Pre-trained transformer-based language models have achieved state-of-the-art accuracy on a wide range of NLP tasks and have shown significant general knowledge capabilities. More recently, new techniques have emerged to improve the reasoning capabilities of LLMs, such as Chain-of-Thought (CoT) prompting and reasoning architectures. CoT is a prompting technique that encourages step by step reasoning within a traditional LLM and has shown a significant improvement in complex reasoning tasks [20]. Reasoning models are specific architectures designed for reasoning often trained with reinforcement learning and using reasoning tokens to invoke an internal CoT-like process [21].
@@ III Problem Formulation
$\tilde{S}=F(S,b,f),\vspace{-10pt}$ (1)
[…]
@@ IV-A Overview
In stage 2, CROP is applied to an unseen design of user query. The unseen design is also processed via LLM-based design analysis and design embedding generation. The generated design embedding is then used to retrieve the parameter guidance of the design with the highest similarity in the database. The retrieved parameter guidance is then utilized by an LLM-based parameter search engine. We detail the embedding-based design retrieval method in subsection IV-C and the LLM-based parameter search engine in subsection IV-D.
@@ IV-D RAG-enhanced LLM Search
The RAG-enhanced LLM search employs a decoder-only LLM as the parameter search engine. During each iteration, CROP constructs a prompt using general information, the parameter names and valid options, data retrieved through the RAG module, and the historical results of previous iterations. This prompt is provided to the model through an API which returns a response that is parsed for the suggested parameter sample. This sample is then validated and evaluated using RTL Architect, and the result is included in the prompts of future iterations. The general information consists of background and instructions. The background explains that the model is tuning EDA flow parameters for PPA objectives. The instructions outline the desired response format, as well as informing the model to be aware of both exploration and exploitation, fostering a more diverse coverage of the search space. We use a chain-of-thought model, as they have been shown to excel at tasks that require advanced scientific reasoning and domain-specific knowledge [26, 21]. This process allows an LLM to act as a fully autonomous design engineer leveraging prior knowledge while thoroughly exploring the search space, freeing human engineers from this laborious task.
@@ V-B Circuit Retrieval Quality
Figure 3: A t-SNE-based visualization of design description embeddings colored by each design’s category, designs from the RTLRewrite benchmark.
@@ V-C Parameter Tuning Performance
Table VI: Comparison of optimization algorithms on Target Core (Power in mW).
@@
In these experiments we report power as the primary metric being optimized however, at the end of this section we present an analysis of the tradeoff between different PPA metrics. Figure 4 compares the performance of CROP against random search, BO and Optuna on Target Core. The results shown in Table VI demonstrate that CROP consistently outperforms all baseline methods across various search budget sizes. Specifically, CROP achieves a clear advantage over the baselines reaching 127.08 mW after only 5 iterations, and converges to a minimal power of 119.98 mW throughout the experiment. This accelerated convergence behavior suggests that the context enrichment from retrieving similar circuits (i.e., Prior Core) provides valuable guidance to effectively constrain the parameter search process. On the other hand, the lowest total power achieved by all other baselines after 75 iterations is 133.17 mW, which means that CROP achieved a 9.9% reduction of the total power consumption compared to all baseline approaches. This substantial improvement highlights the efficacy of knowledge transfer between similar designs implemented in different technology nodes. Notably, while traditional optimization methods continue to make incremental improvements throughout their search trajectories, they fail to identify the high-quality parameter regions that CROP discovers almost immediately through its retrieval-augmented approach, demonstrating the limitations of design-agnostic search strategies for modern VLSI designs.
@@
V-D RAG Benefit Analysis
In order to evaluate the efficacy of the RAG component within CROP we compare CROP to a naive LLM-based search using the same model prompted without the information from the design retrieval component. We ran multiple iterations of both searches, which are shown in Figure 5. These results show a 7.5% and 7.8% reduction in power usage after 5 and 40 iterations, respectively, when including the RAG information in the search prompting. The performance of the naive LLM is comparable to those of the baseline methods shown in Figure 4 showing that LLM guided search on its own is a valid method for parameter tuning. This relies on the LLM’s ability to intuitively understand the impact of parameters, tracking how they effect the result over time using only the knowledge encoded in the LLM’s training. However, the RAG-aware LLM significantly outperforms any other method. The benefit of this RAG information is shown in how quickly it is able to find significantly better values than the other approaches, in some cases on the first iteration, as shown in the CROP_2 run. These results highlight both the efficacy of the RAG component in CROP’s search and the ability of CROP to save significant amounts of time and computational resources.
Figure 5: Two iterations each of CROP and CROP without RAG information to show the advantage of RAG on Target Core.
@@ V-F Runtime Analysis
Table VII: Runtime Breakdown by Framework Component for Optimizers (Target Core).
@@
As the overall goal of this work is to reduce the time parameter tuning takes, it is important to evaluate the runtime of our approach. Table VII shows the runtime breakdown for CROP as well as BO and Optuna. The runtime analysis of each of these methods reveals that running the EDA flow dominates the total runtime of the experiments. CROP has additional initialization overhead from the RTL to embedding extraction and similarity-base retrieval mechanism, however, this overhead’s impact is significantly reduced as more iterations are run. During each iteration CROP’s parameter selection stage is mostly spent waiting on the inference of the reasoning model while Optuna and BO’s are dominated by logging and file clean up with the actual next parameter calculation taking a small fraction of that time. Additionally, the baselines have negligible initialization overhead; however, the running of the EDA flow is no faster. The amount to which the EDA flow dominates the runtime of all approaches leads to very marginal differences between them. Automated approaches such as these will always be faster per iteration than those involving a human expert.
@@ VI Discussion
This method proves particularly valuable for industrial-scale designs where direct embedding of complete RTL code would exceed even the expanded context windows of modern LLMs (e.g., 128K tokens for GPT-4o in our experiment setting). Our experiments confirm that these derived embeddings successfully capture semantic similarities between functionally related circuits, as evidenced by the clustering patterns in Figure 3, while avoiding the prohibitive computational costs of processing entire RTL repositories directly.
@@
The experimental evaluations of CROP on industrial design show that it is a promising direction for improving automated flow tuning. The inherent database reliance of CROP is both a strength and a weakness; it requires similar designs to be present in the database in order to achieve best results but as more designs are ran through the framework they are added to the database expanding the capability of CROP. Furthermore, CROP is directly compatible with existing approaches that use jump-start or early-stopping methods. CROP is an EDA tool agnostic method that allows it to be used with any tool-chain, existing methods such as [4, 10] implement methods to reduce the runtime of the EDA tools. As shown in our run-time analysis, this has the potential to even further improve the overall efficiency of CROP. Additionally, CROP can be combined with search parallelization methods such as the creation of separate trust regions to search within. We hypothesize that this is a potentially highly beneficial direction of improvement as the results retrieved from the database can be used to reduce the number of initial iterations required to establish trust regions which can then be searched in parallel using multiple instances of LLM guided search.
@@ Acknowledgment
This work is supported by NSF 2106828 and Synopsys Gift.

### v2

@@ I Introduction
To address this challenge, a promising research direction leverages circuit characteristics to enhance parameter tuning. This circuit-aware approach incorporates prior knowledge of design structure and functionality to more efficiently constrain the optimization process. VLSI design flows implementing such circuit-aware parameter tuning have demonstrated significant improvements in search efficiency. [13] used knowledge provided from circuit designers to guide the genetic algorithm and improve search efficiency. [4] used random embedding to convert the parameters to a lower-dimensional latent space, which can be more efficiently explored based on design similarity. [10] exploited archival design data to improve the warm-up efficiency of the search engine. Despite these advances, significant challenges remain in effectively transferring knowledge in a way that better captures the correlation between pre-existing archival design data and the design currently being optimized. This is due to insufficient research on the effective encoding of the semantic representations of circuit designs. An approach that can understand circuit semantics and leverage knowledge from similar designs could dramatically reduce the number of iterations required to find optimal parameters.
Building on recent advances in LLMs and their demonstrated effectiveness in EDA and chip design tasks [14, 15, 16], we investigate the potential of LLMs to enable fully-automated circuit-aware parameter search engines. Unlike traditional block-box optimization tools, LLMs can be conditioned using rich descriptive prompts that incorporate the QoR target, the functionality of each tool option, and even high-level descriptions of the circuit under optimization. This allows LLMs to emulate the intuition and practice of experienced engineers, making context-aware adjustments to EDA tool configurations, design constraints, and metric-oriented strategies. Furthermore, RAG emerges as a systematic approach that augments general-purpose LLMs with a domain-specific knowledge database, enabling more informed parameter optimization decisions based on historical design data and established optimization patterns. The integration of RAG with LLMs creates a parameter tuning framework that leverages contextual reasoning with circuit-specific prior knowledge, enabling parameter optimization that adapts to unique design characteristics across diverse technology nodes.
@@ II Preliminaries
Figure 1: Overview of CROP. In stage 1, CROP constructs a design database. CROP first performs the two-step LLM-based design analysis to produce the design embedding. CROP selects the top-$k$ parameter sets and their QoR results from pre-existing optimization results to compile the parameter guidance for the circuits in the database. In stage 2, CROP conducts retrieval-augmented parameter search. First, CROP performs LLM-based analysis on the design of user query and retrieves the parameter guidance of the design from the database with highest similarity based on maximum inner product search (MIPS). The retrieved parameter guidance is then used to constrain the LLM-based parameter search engine that optimizes the QoR for the input design.
@@ II-A Transformer-based Language Models
The advent of transformer-based language models has revolutionized the field of natural language processing (NLP) resulting in significant advancements in numerous tasks such as question answering, text generation, and machine translation. Introduced in 2017 by [17] the transformer architecture replaced recurrence with multihead self-attention, a mechanism that allowed more efficient training and better performance in sequence modeling tasks. Subsequent developments of larger models such as GPT3 [18] and LLaMA3 [19], have shown strong zero-shot and few-shot generalization skills when trained on massive corpora. These models are typically pre-trained for next-token generation and later finetuned for downstream tasks such as instruction following or question answering. Pre-trained transformer-based language models have achieved state-of-the-art accuracy on a wide range of NLP tasks and have shown significant general knowledge capabilities. More recently, new techniques have emerged to improve the reasoning capabilities of LLMs, such as Chain-of-Thought (CoT) prompting and reasoning architectures. CoT is a prompting technique that encourages step by step reasoning within a traditional LLM and has shown a significant improvement in complex reasoning tasks [20, 21]. Reasoning models are specific architectures designed for reasoning often trained with reinforcement learning and using reasoning tokens to invoke an internal CoT-like process [22].
@@ III Problem Formulation
$\tilde{S}=F(S,b,f),\vskip-10.0pt$ (1)
$F^{*}=\underset{F}{\arg\min}(\min O[\tilde{S}]-\min O[S]).$ (2)
@@ IV-A Overview
In stage 2, CROP is applied to an input design of user query. The input design is also processed via LLM-based design analysis and design embedding generation. The generated design embedding is then used to retrieve the parameter guidance of the design with the highest similarity in the database. The retrieved parameter guidance is then utilized by an LLM-based parameter search engine. We detail the embedding-based design retrieval method in subsection IV-C and the LLM-based parameter search engine in subsection IV-D.
@@ IV-D RAG-enhanced LLM Search
The RAG-enhanced LLM search employs a decoder-only LLM as the parameter search engine. During each iteration, CROP constructs a prompt using general information, the parameter names and valid options, data retrieved through the RAG module, and the historical results of previous iterations. This prompt is provided to the model through an API which returns a response that is parsed for the suggested parameter sample. This sample is then validated and evaluated using RTL Architect, and the result is included in the prompts of future iterations. The general information consists of background and instructions. The background explains that the model is impersonating an experienced EDA engineer and tuning EDA flow parameters for QoR objectives. The instructions outline the desired response format, as well as informing the model to be aware of both exploration and exploitation. We use a CoT model, as they have been shown to excel at tasks that require advanced scientific reasoning and domain-specific knowledge [27, 22]. Additionally, following [21], we construct our prompts with clear instructions which makes the CoT model adhere to a consistent output format. This process allows an LLM to act as a fully autonomous design engineer leveraging prior knowledge while thoroughly exploring the search space, freeing human engineers from this laborious task.
@@ V-B Circuit Retrieval Quality
Figure 3: A t-SNE-based visualization of design description embeddings colored by each design’s category, designs from the RTLRewriter benchmark.
@@ V-C Parameter Tuning Performance
Table VI: Comparison of Target Core’s total power consumption (mW) using different parameter optimization algorithms.
@@
In these experiments we report power as the primary metric being optimized however, at the end of this section we present an analysis of the tradeoff between different PPA metrics. Figure 5 compares the performance of CROP against random search, BO and Optuna on Target Core. The results shown in Table VI demonstrate that CROP consistently outperforms all baseline methods across various search budget sizes. Specifically, CROP achieves a clear advantage over the baselines reaching 127.08 mW after only 5 iterations, and converges to a minimal power of 119.98 mW throughout the experiment. This accelerated convergence behavior suggests that the context enrichment from retrieving similar circuits (i.e., Prior Core) provides valuable guidance to effectively constrain the parameter search process. On the other hand, the lowest total power achieved by all other baselines after 75 iterations is 133.17 mW, which means that CROP achieved a 9.9% reduction of the total power consumption compared to all baseline approaches. This substantial improvement highlights the efficacy of knowledge transfer between similar designs implemented in different technology nodes, as well as the LLMs ability to guide efficient search. Notably, while traditional optimization methods continue to make incremental improvements throughout their search trajectories, they fail to identify the high-quality parameter regions that CROP discovers almost immediately through its retrieval-augmented approach, demonstrating the limitations of design-agnostic search strategies for modern VLSI designs.
Figure 4: Comparison of CROP and naive LLM search (without RAG-based parameter guidance retrieval) across two independent runs each.
V-D RAG Benefit Analysis
In order to evaluate the efficacy of the RAG component within CROP we compare CROP to a naive LLM-based search using the same model prompted without the information from the design retrieval component. We ran multiple iterations of both searches, which are shown in Figure 4. Including RAG information in the search prompting resulted in a 6.59% and 8.66% reduction in power usage after iterations 5 and 40, respectively, compared to the best values from naive LLM search traces. The performance of the naive LLM is comparable to those of the baseline methods shown in Figure 5 showing that LLM guided search on its own is a valid method for parameter tuning. This relies on the LLM’s ability to intuitively understand the impact of parameters, tracking how they effect the result over time using only the knowledge encoded in the LLM’s training. However, the RAG-aware LLM significantly outperforms any other method. In CROP, RAG enables faster convergence to optimal values, achieving superior results with fewer iterations across all trials. These results highlight both the efficacy of the RAG component in CROP’s search and the ability of CROP to save significant amounts of time and computational resources.
@@ V-F Runtime Analysis
Table VII: Runtime breakdown by framework component for optimizers on Target Core.
@@
As the overall goal of this work is to reduce the time parameter tuning takes, it is important to evaluate the runtime of our approach. Table VII shows the runtime breakdown for CROP as well as BO and Optuna. The runtime analysis of each of these methods reveals that running the EDA flow dominates the total runtime of the experiments. The runtime analysis reveals that running the EDA flow dominates the total runtime across all methods, leading to relatively marginal differences between approaches.
However, several distinctions emerge in the initialization and iteration phases. CROP incurs additional initialization overhead from the RTL to embedding extraction and similarity-base retrieval mechanism, while the baselines have negligible initialization costs. Nevertheless, CROP’s initialization overhead is amortized across subsequent iterations, reducing its relative impact as more iterations are run. The methods also differ in how they spend time during each iteration. CROP’s parameter selection stage is mostly spent waiting on the inference of the reasoning model, whereas Optuna and BO’s are dominated by logging and file cleanup, with the actual next parameter calculation requiring a small fraction of that time.
While automated approaches will always outperform human experts on a per-iteration basis, CROP’s key advantage lies in solution quality rather than speed per iteration. CROP’s ability to find notably better solutions within significantly fewer iterations allows for the potential of early stopping leading to a massive save in overall runtime compared to BO and Optuna.
@@ VI Discussion
This method proves particularly valuable for industrial-scale designs where direct embedding of the complete RTL code could easily exceed even the expanded context windows of modern LLMs (e.g., 128K tokens for GPT-4o in our experimental setting). Our experiments confirm that these derived embeddings successfully capture semantic similarities between functionally related circuits, as evidenced by the clustering patterns in Figure 3, while avoiding the prohibitive computational costs of processing entire RTL repositories directly.
@@
The experimental evaluations of CROP on industrial designs show that it is a promising direction for improving automated flow tuning. We identify the following areas for future improvement: 1) The inherent database reliance of CROP presents both strengths and challenges; while it demonstrates the strength of achieving better QoR results in significantly fewer iterations as shown in our evaluations, the challenge emerges as the database expands and may require a more sophisticated method for design retrieval that maintains accuracy and efficiency. 2) This retrieval-driven approach is complemented by CROP’s tool-agnostic nature, allowing seamless integration with existing EDA toolchains and compatibility with approaches that use jump-start or early-stopping methods. Methods such as REMOTune [4] and FlowTuner [10] that reduce EDA tool runtime can be directly incorporated, potentially further improving CROP’s overall efficiency as demonstrated in our runtime analysis. 3) Significant performance gains can be achieved through parallelization, as CROP can be combined with search parallelization methods such as separate trust regions. This represents a highly beneficial improvement direction since database results can reduce initial iterations required to establish trust regions, which can then be searched in parallel using multiple LLM-guided search instances. 4) CROP’s model-agnostic design enables improvements through enhanced LLMs, whether by adopting newer and more capable models, implementing fine-tuned domain-specific models, or applying training-free decoding techniques that reduce hallucination [34, 35].
@@ Acknowledgment
This work is in part supported by NSF 2106828 and Synopsys gift.